OpenStudy (anonymous):

How are the formulas for the circumference of a circle and area of a circle derived?

OpenStudy (anonymous):

ecall that the definition of pi (π) is the circumference c of any circle divided by its diameter d. Put as an equation, pi is defined as Rearranging this to solve for c we get The diameter of a circle is twice its radius, so substituting 2r for d If you know the area Recall that the area a of a circle is given by Solving this for r The circumference c of a circle is Substituting r Squaring both sides to lose the radical Solving for c:

OpenStudy (anonymous):

recall*

OpenStudy (here_to_help15):

well circle is actually a sector with an angle of 2pi radians b/w it's arms i know it sounds freaky but it's actually that the sector tends towards the circle and the length of the arc formula derivation can be found on sum website the formula is l = (angle in radians)(distance b/w the points of the locus and centre of rotation i.e. radius) so l = (angle)r in case of a circle angle = 2pi radians so l = 2*pi*r here the arc would have tended to circumference as sector tends towards circle so C = 2*pi*r units

OpenStudy (anonymous):

That is the correct answer yw

OpenStudy (anonymous):

I know that the formula to find the circumference of a circle is C=3.14d I know that the area of the circle formula is A=3.14r2

OpenStudy (anonymous):

ok well then this should be a piece of cake

OpenStudy (anonymous):

I am confused on what the answer is. How ould i answer this question simply? @ArmyBoy97

OpenStudy (here_to_help15):

ummm let me see.

OpenStudy (anonymous):

derived means to take or get (something) from something else i think

OpenStudy (anonymous):

sin(theta) = a / c csc(theta) = 1 / sin(theta) = c / a cos(theta) = b / c sec(theta) = 1 / cos(theta) = c / b tan(theta) = sin(theta) / cos(theta) = a / b cot(theta) = 1/ tan(theta) = b / a

OpenStudy (anonymous):

my bad wrong question

OpenStudy (here_to_help15):

that is correct @EmmaDrew94

OpenStudy (anonymous):

lol one sec

OpenStudy (anonymous):

i am helping two people at once

OpenStudy (anonymous):

@ArmyBoy97 ok your just confusing me that isnt helping me at all. What is the answer thats not where the formulas derive from

OpenStudy (anonymous):

im soryy im multitasking righht now

OpenStudy (here_to_help15):

Ok wait a sec i'll try:)

OpenStudy (anonymous):

@Here_to_Help15 he will help you im really busy

OpenStudy (nikato):

well im not sure if this is what your looking for. but for circuimference, if you take the diameter and wrap it around the circle, you would need about 3 and a little more of the diameter to completely wrap around the circle

OpenStudy (anonymous):

@nikato deserves a medal

OpenStudy (anonymous):

go away lol i thought you were busy @ArmyBoy97

OpenStudy (anonymous):

ok whatever i was just trying to help not my fault your not good at math?

OpenStudy (anonymous):

Thank you @nikato

OpenStudy (anonymous):

you shouldnt be rude to people who try to help you

OpenStudy (anonymous):

then no one will want to help you